OpenStudy (cheska_p):

I'm a visual learning, can some one please work this out so I can see how you get the answer? Physics exam in two days! The velocity v(t) of a particle as a function of time is given by v(t) = (2.3 m/s) + (4.1 m/s2)t - (6.2 m/s3)t2. What is the average acceleration of the particle between t = 1.0 s and t = 2.0 s (up to two significant figures) ?

OpenStudy (anonymous):

\[Average acceleration=\frac{ Final velocity-initial velocity }{ time 2-time1 }\]

OpenStudy (cheska_p):

The problem I'm having is that I don't know what to do with that long particle function. I tried differentiating it but I think I'm doing it wrong

OpenStudy (anonymous):

Just plug in the first time period to get your initial velocity.

OpenStudy (anonymous):

\[2.3+4.1t-6.2t^2\] You can shorten it like this.

OpenStudy (cheska_p):

I figured it out! Thank you. I do have another question though, why is it that the units (m/s, m/s^2, m/s^3) don't matter?

OpenStudy (anonymous):

It does matter. Better to get it out of the way to get initial and final velocity.

OpenStudy (anonymous):

Are you talking about the final answer? It should be in m/s^2 which is \[m/s^2\]

OpenStudy (cheska_p):

No I was referring to the particle function, you really helped me understand thank you!

OpenStudy (anonymous):

It actually does

OpenStudy (anonymous):

\[m/s+\frac{ m \times s }{ s^2 }+\frac{ m \times s^2 }{ s^3 }=m/s+m/s+m/s\]

OpenStudy (anonymous):

Time is in seconds.